An automobile engine provides 532 Joules of work to push the pistons.
In this process the internal energy changes by -2872 Joules.

What is q for the engine? This represents the amount of heat that must be carried away by the cooling system.
q = _______ Joules
1
Expert's answer
2019-04-01T07:48:58-0400

Solution.

Q = U + A

Q<0, because heat that must be removed by the cooling system.

Q = -2872 + 532 = -2340 J

Answer:

Q = -2340 J